OpenStudy (anonymous):

A balloon rises at a rate of 4 meters per second from a point on the ground 40 meters from an observer. What is the rate of change of the angle of elevation (in rad/sec) when the balloon is 40 meters above the ground?

OpenStudy (anonymous):



OpenStudy (anonymous):

that's the equation relating y and theta...

OpenStudy (anonymous):

What does that equation look like?

OpenStudy (anonymous):

differentiate with respect to time, solve for dtheta/dt

OpenStudy (anonymous):

\[\tan \theta =\frac{ y }{ 40}\]

OpenStudy (anonymous):

differentiate and solve for d theta / dt.. plug in values for y, y' and theta...

OpenStudy (anonymous):

think you got it now?

OpenStudy (anonymous):

Can you please show me the set up of this problem because I do not understand.

OpenStudy (anonymous):

you're doing related rates, you've done a few problems where you differentiate with respect to time? using the chain rule?

OpenStudy (anonymous):

should be familiar...

OpenStudy (anonymous):

Yes I have but I've never taken Trig...

OpenStudy (anonymous):

eg. if x and y are functions of time... then ( x^2 +y^2 ) ' = 2x*dx/dt + 2y*dy/dt etc

OpenStudy (anonymous):

so... what's your question about trig.

OpenStudy (anonymous):

tangent is opp/adj

OpenStudy (anonymous):

Right. I don't understand where do I go from tan delta = y / 40

OpenStudy (anonymous):

don't now how you're in calc. without having taken trig. ever, but I guess there are alternate universes where anything can happen

OpenStudy (anonymous):

tan(theta) theta is just the variable used for angles..

OpenStudy (anonymous):

it's like x

OpenStudy (anonymous):

you've done derivatives of trig functions?

OpenStudy (anonymous):

derivative of tan is?

OpenStudy (anonymous):

Sec squared

OpenStudy (anonymous):

yes... and using the chain rule too...: sec^2 (theta) * d (theta) /dt

OpenStudy (anonymous):

that's the left side...

OpenStudy (anonymous):

right side is just dy/dt * 1/40

OpenStudy (anonymous):

\[\sec ^{2} \theta \frac{ d \theta }{ dt} = \frac{ dy }{ dt }\frac{ 1 }{40 }\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Okay. I understand that. I'm solving for dy/dt right?

OpenStudy (anonymous):

naw it asks for d theta /dt

OpenStudy (anonymous):

"What is the rate of change of the angle of elevation ..."

OpenStudy (anonymous):

find theta.. divide both sides by sec^2(theta)

OpenStudy (anonymous):

plug in dy/dt and that's it...

OpenStudy (anonymous):

theta you find from "What is the rate of change of the angle of elevation (in rad/sec) when the balloon is 40 meters above the ground?" so when y = 40m theta = ..?

OpenStudy (anonymous):



OpenStudy (anonymous):

4?

OpenStudy (anonymous):

45 deg

OpenStudy (anonymous):

you should get in the habit of using radians however... so happens that it doesn't matter in this case... but it usually matters if you're doing calculus with angles... so pi/4 is the angle in radians

OpenStudy (anonymous):

switch to radian mode on your calculator and find cos(pi/4) then square it..

OpenStudy (anonymous):

k got it?

OpenStudy (anonymous):

I got .5

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

.5* dy/dt * 1/40 is d(theta) / dt

OpenStudy (anonymous):

Is that my answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Thank You So Much! Sorry , for my lack of Trig knowledge making your job harder.

OpenStudy (anonymous):

no problem... is it clearer now?

OpenStudy (anonymous):

Yes. Very clear.

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

(: